OpenStudy (anonymous):

lim t→1 e^−t(cos(πt)+e^t)/3. ?

OpenStudy (anonymous):

\[\lim_{t \rightarrow 1}\frac{ e⁻^t (\cos(\Pi t))+e^t}{ 3 }\]

OpenStudy (anonymous):

the answer is \[\frac{ 1 }{ 3 }-\frac{ e⁻¹ }{ 3 }\] but I don't understand how

OpenStudy (harsimran_hs4):

it`s a direct question all you need to do is to put value of t =1 and get the answer

OpenStudy (anonymous):

but if I put value t=1 I get decimal numbers?

OpenStudy (anonymous):

Like, maybe my problem is that I don't understand what cos(pi) is?

OpenStudy (harsimran_hs4):

cos(pi) = cos(180 degree) = -1